OpenStudy (anonymous):

If the degree measure of an angle is 40° less than 3 times the degree measure of its supplementary angle, what is the degree measure of the angle? A. 50° B. 55° C. 125° D. 130°

OpenStudy (cwrw238):

x = 3 (180 - x) - 40 solve for x

OpenStudy (cwrw238):

can you solve this equation?

OpenStudy (anonymous):

so the equation would x = 540 - 3x - 40

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

do i add 40 to 540

OpenStudy (cwrw238):

no subtract 540 - 40 = 500 and equation becomes x = -3x + 500

OpenStudy (cwrw238):

now we aim to get all x terms on the left: add 3x to both sides of = sign: x + 3x = -3x + 3x + 500 4x = 500 can you continue?

OpenStudy (anonymous):

divide by 4 and you get 125

OpenStudy (cwrw238):

right...

OpenStudy (cwrw238):

C is correct - good job